what are the intercepts to the equation


4x^2+16y^2=64

Answer by rapaljer(4671)

To find the x intercepts, let y =0.
If y=0, then 4x^2=64
x^2 = 64/4 =16
x= 4 or x= -4

To find the y intercepts, let x=0.
If x=0, then 16y^2 = 64
y^2 = 64/16=4
y= 2 or y=-2.

R^2

Dr. Robert J. Rapalje, Retired
Seminole State College of Florida
Altamonte Springs Campus